# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ 5.4 Randomized Lower-Bound for $r$: Inner Radius of a Ridge Projected onto $(d-1)$-Dimensional Space
5.5 Combining Together and Proof of Lemma 33
@@
In this section, we combine the deterministic argument in Lemma 36, and the randomized arguments in Lemma 38 and Lemma 42. We can finally show the main technical lemma (Lemma 33).
Proof of Lemma 33.
@@
Without loss of generality, let $I=[d]$ and write $E=E_{I}$. Suppose $p^{\prime}=A_{J}=\operatorname{conv}(a_{j}:j\in J)\cap W$ is the next vertex after the edge $F_{I}\cap W$. Here $J\in\binom{[n]}{d-1}$ and $R_{J}$ is the $(d-2)$-dimensional ridge. With probability $1$, the polytope $\operatorname{conv}(a_{1},\dots,a_{n})$ is non-degenerate and $W\cap R^{\prime}$ is a single point for any ridge $R^{\prime}$ of $\operatorname{conv}(a_{1},\dots,a_{n})$ that intersects with $W$. We will show that conditional on $E$, each of the following conditions in the deterministic argument (Lemma 36) is satisfied with good probability:
@@
    (Lower bound of $\delta$) $\min_{k\in[n]\backslash I}\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})\geq\Omega(\frac{1}{Ld\log n})$;
@@
Note for the last point that Lemma 36 only requires this for the set $J$ which indexes the second vertex of $F_{I}\cap W$ in clockwise direction, but we prove it for both of the sets $J$ for which $R_{J}\cap W\neq\emptyset$.
First, we write $D$ as the event that $\forall i,j\in[n]$ for which $\|a_{i}-a_{j}\|\leq 3$. From Lemma 16, for any $\sigma\leq\frac{1}{8\sqrt{d\log n}}$, with probability at least $1-\binom{n}{d}^{-1}$, we have $\max_{i\in[n]}\|\hat{a}_{i}\|\leq 1+4\sigma\sqrt{d\log n}\leq\frac{3}{2}$, i.e. $\Pr[D]\geq 1-\binom{n}{d}^{-1}$. Using the assumption that $\Pr[E_{I}]\geq 10\binom{n}{d}^{-1}$, we have
@@
Next, we consider $\delta:=\operatorname{dist}(\operatorname{affhull}(a_{1},\dots,a_{d}),\{a_{d+1},\dots,a_{n}\})$. Using Lemma 38, we have $\Pr[\delta\geq\frac{1}{10e^{3}Ld\log n}\mid E]\geq 0.72$.
Finally, we consider $r:=\max_{J}\operatorname{dist}(\operatorname{affhull}(a_{1},\ldots,a_{d})\cap W,\partial R_{J})$ subject to all $J\in\binom{I}{d-1}$ such that $A_{J}$ happens (in other words, $R_{J}=\operatorname{conv}(a_{j}:j\in J)$ is a ridge of $F_{I}$ such that $R_{J}\cap W\neq\emptyset$). By union bound,
@@
From Lemma 42, for each $J\in\binom{I}{d-1}$, we know that
@@
$\geq$ $1-0.1\cdot 2-2\cdot\Pr[\neg D\mid E]\geq 0.6$
Therefore by union bound, the three conditions hold with probability at least $1-(1-0.9)-(1-0.72)-(1-0.6)\geq 0.1$, and the lemma directly follows from Lemma 36. ∎
@@ 6 Smoothed Complexity Lower Bound
In this section, we present the lower bound of the smoothed complexity by studying the intersection between the smoothed dual polytope $Q=\operatorname{conv}(a_{1},\ldots,a_{n})\subseteq\mathbb{R}^{d}$ (where each $a_{i}$ is under Gaussian perturbation), and the two-dimensional shadow plane $W\subseteq\mathbb{R}^{d}$. Our main result is as follows:
Theorem 49.
For any $d>5,n=4d-13$, there exists a two-dimensional linear subspace $W\subseteq\mathbb{R}^{d}$ and vectors $\bar{a}_{1},\ldots,\bar{a}_{n}\in\mathbb{R}^{d}$, $\max_{i\in n}\|\bar{a}_{i}\|_{1}\leq 1$ such that the following holds. Let $a_{1},\ldots,a_{n}$ be independent Gaussian random variables where each $a_{i}\sim\mathcal{N}_{d}(\bar{a}_{i},\sigma^{2}I),\sigma\leq\frac{1}{7200d\sqrt{\log n}}$. Then with probability at least $1-\binom{n}{d}^{-1}$, we have
@@
Theorem 49 is the direct consequence of the next theorem, which is a lower bound for adversarial perturbations of bounded magnitude:
Theorem 50.
For any $d>5,n=4d-13$, there exists a two-dimensional linear subspace $W\subseteq\mathbb{R}^{d}$ and vectors $\bar{a}_{1},\ldots,\bar{a}_{n}\in\mathbb{R}^{d}$, $\max_{i\in n}\|\bar{a}_{i}\|_{1}\leq 1$ such that the following holds. For any $\varepsilon<\frac{1}{180}$, if $a_{1},\ldots,a_{n}\in\mathbb{R}^{d}$ satisfy $\|a_{i}-\bar{a}_{i}\|_{1}\leq\varepsilon$ for all $i\in[n]$ then we have
@@
The rest of this section is organized as follows: In Section 6.1, we construct the polytope $P$ by a set of constraints, and the two-dimensional shadow plane $W$. An informal intuition behind these inequalities is described in Section 6.2. In Section 6.3, we show that $\pi_{W}(P)$ approximates the unit disk $\mathbb{B}_{2}^{2}$. In Section 6.4, we analyze the largest $\ell_{\infty}$-ball contained in $P$ and the smallest $\ell_{\infty}$-ball containing $P$. Section 6.5 investigates the polar polytope $Q=(P-x)^{\circ}$ of a shift of $P$ and derives bounds on the radius of its largest contained $\ell_{1}$-ball and smallest containing $\ell_{1}$-ball. Finally, Section 6.6 shows that the bounded ratio between these radii imply that any perturbation $\tilde{Q}$ still has $\tilde{Q}\cap W$ approximate the unit disk $\mathbb{B}_{2}^{2}$ well and uses this fact to prove Theorem 49.
6.1 Construction of the Primal Polytope
In this subsection, we first construct the primal polytope and the two-dimensional plane $W$. For $k\in\mathbb{N}$, we construct a $(k+5)$-dimensional polytope. We will use the following vectors in the definition:
@@ 6.1 Construction of the Primal Polytope
With these definitions in mind, let $P^{\prime}\subseteq\mathbb{R}^{3k+5}$ denote the set of points $(x,y,p_{0},\dots,p_{k},t,s)$, where $x,y\in\mathbb{R},p_{0},p_{1},\ldots,p_{k}\in\mathbb{R}^{2},t\in\mathbb{R}^{k},s\in\mathbb{R}$, satisfying the following system of linear inequalities:
@@
$\mathbf{0}_{k}\preceq t$ $\preceq\mathbf{1}_{k}$ (26)
@@
The plane $W$ of interest is that spanned by the unit vectors in the $x$ and $y$ directions.
An illustration of the vertices of the projected polytope $\pi_{W}(P)$ can be found in Figure 3 for $k=4$. Note that the figure appears to depict a regular polygon with $2^{k+1}$ vertices.
Figure 3: Vertices of the projected primal polytope $\pi_{W}(P)$ (see (28)) without perturbation for $k=4$.
6.2 Intuition of the Construction
To explain the intuition behind the equations (22 - 27), let us consider the simpler system of inequalities in variables $x,y\in\mathbb{R},r_{0},\dots,r_{k}\in\mathbb{R}^{2}$:
$e_{1}^{\top}r_{0}$ $\geq|x|,e_{2}^{\top}r_{0}\geq|y|$ (29)
$w_{i}^{\top}r_{i}$ $\geq w_{i}^{\top}r_{i-1},~{}\forall i\in[k]$ (30)
$v_{i}^{\top}r_{i}$ $\geq|v_{i}^{\top}r_{i-1}|,~{}\forall i\in[k]$ (31)
$e_{1}^{\top}r_{k}$ $\leq 1$ (32)
$e_{2}^{\top}r_{i}$ $\geq 0,~{}\forall i\in[k]$ (33)
Let $R\subseteq\mathbb{R}^{2k+4}$ denote the set of vectors $(x,y,r_{0},\dots,r_{k})$ satisfying the above inequalities. For each $i=0,\dots,k$, write $R_{i}=\{r_{i}:\exists x,y,r_{0},\dots,r_{i-1},r_{i+1},\dots,r_{k}\text{ s.t. }(x,y,r_{0},\dots,r_{k})\in R\}\subseteq\mathbb{R}^{2}$ for the projections of $R$ onto the two-dimensional coordinate subspace of $r_{i}$. Also, let $W$ be the two-dimensional plane spanned by the $x$ and $y$ directions, so that $\pi_{W}(R)=\{(x,y):\exists r_{0}\in R_{0},r_{0}\geq(|x|,|y|)^{\top}\}$. The vertices of these sets are depicted in Figure 4 for $k=2$.
For these sets, we have the following observations. We will skip the proof since they only give illustrations of our analysis and will not be functional to the proof of Theorem 49.
  1. 1.
    The set $R_{k}$ can be described by the inequalities $e_{2}^{\top}r_{k}\geq 0,e_{1}^{\top}r_{k}\leq 1,v_{k}^{\top}r_{k}\geq 0$. These inequalities describe a small slice of the regular $2^{k+1}$-gon.
  2. 2.
    For each $i=0,\dots,k-1$, the set $R_{i}$ is obtained by taking the union of $R_{i+1}$ and its mirror image in the line spanned by $w_{i}$.
  3. 3.
    For each $i=1,\dots,k$, the set $R_{i}$ can be described as the restriction of a regular $2^{k+1}$-gon intersected with the set $\{r_{i}:e_{2}^{\top}r_{i}\geq 0,v_{i}^{\top}r_{i}\geq 0\}$.
  4. 4.
    The set $R_{0}$ can be described as a regular $2^{k+1}$-gon intersected with the non-negative orthant.
  5. 5.
    The set $\pi_{W}(R)=\{(x,y):\exists r_{0}\in R_{0},r_{0}\geq(|x|,|y|)^{\top}\}$ is a regular $2^{k+1}$-gon.
More careful inspection allows for the following additional observations, which we also state without proof:
  1. 1.
    Turning (30) from an inequality to an equality constraint does not change any of the sets $R_{0},\dots,R_{k}$.
  2. 2.
    Removing the constraint (33) does not change $R_{0}$.
  3. 3.
    Adding upper bounds $v_{i}^{\top}r_{i}\leq 2$ does not change $R_{0}$.
Each of the above-mentioned changes either serves to increase the radius of the largest ball (in the affine hull of $R$) contained in the relative interior of $R$ or to decrease the radius of the smallest ball containing $R$. The addition of the variable $s$ in the construction of $P$ serves to further increase the radius of the largest ball contained in $P$ without increasing the radius of the smallest ball containing $P$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Vertices of $\pi_{W}(R),R_{0},R_{1},R_{2}$ for $k=2$.
6.3 Projected Primal Polytope Approximates Two-Dimensional Unit Disk

### v3

@@ 5.4 Randomized Lower-Bound for $r$: Inner Radius of a Ridge Projected onto (d1)(d-1)-Dimensional Subspace
Now we show Lemma 5.20.
Proof 5.27 (Proof of Lemma 5.20).
In the following arguments, we condition on $\ell\cap\operatorname{conv}(b_{i}:i\in[d-1])\neq\emptyset$. Without loss of generality, set i=d1i=d-1 and assume that $\ell$ is a linear subspace, i.e., $\pi_{w^{\perp}}(\ell)=0$.
We start with a coordinate transformation. Let $\phi\in w^{\perp}\cap\mathbb{S}^{d-2}$ denote the unit vector satisfying $\phi^{\top}b_{1}=\phi^{\top}b_{j}>0$ for all $j=1,\dots,d-1$. Note that $\phi$ is uniquely defined almost surely: $w^{\perp}$ is a (d2)(d-2)-dimensional linear space and we impose (d3)(d-3) linear constraints $\{\phi^{\top}b_{1}=\phi^{\top}b_{j},\forall j\in[d-2]\}$. Almost surely, these give a one-dimensional linear subspace which, after adding the unit norm and $b_{1}^{\top}\phi>0$ constraint, leaves a unique choice of $\phi$.
Now define $h\in\mathbb{R}$ by $h=\phi^{\top}b_{1}$ and define $\alpha\in\mathbb{R}$ by $\alpha h=-\phi^{\top}b_{d-1}$. Since $0\in\operatorname{conv}(\pi_{w^{\perp}}(b_{i}):i\in[d-1])$ but $\phi^{\top}b_{i}>0$ for all $i\in[d-2]$, we must have $\alpha\geq 0$ for otherwise $\phi$ would separate $\operatorname{conv}(\pi_{w^{\perp}}(b_{i}):i\in[d-1])$ from 0. Again from almost-sure non-degeneracy we get $\alpha>0$ and $h\neq 0$. We define the following coordinate transformation:
$b_{j}$ $=h\phi+c_{j},\quad\forall j\in[d-2]$
$b_{d-1}$ $=-\alpha h\phi+c_{d-1}$
where for each $j\in[d-1]$, $c_{j}\in\phi^{\perp}\cap\operatorname{affhull}(b_{1},\ldots,b_{d-1})$ has (d3)(d-3) degrees of freedom. From here on out, we consider the vertices $(b_{1},\ldots,b_{d-1})$ to be a function of $(h,\alpha,\phi,c_{1},\dots,c_{d-1})$. Again by Lemma 2.13, the induced joint probability density on $(h,\alpha,\phi,c_{1},\dots,c_{d-1},b_{d})$, is proportional to
Using the principle of deferred decision, fix the exact values of $(\alpha,\phi,c_{1},\dots,c_{d-1},b_{d})$. When this is the case we find that
$\propto$~ $\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\dots,b_{d}))\cdot\operatorname{vol}_{d-3}(\operatorname{conv}(\pi_{w^{\perp}}(c_{1}),\dots,\pi_{w^{\perp}}(c_{d-2})))\cdot\prod_{j=1}^{d}\bar{\mu}_{j}(b_{j}).$
To see why $\operatorname{vol}_{d-3}(\operatorname{conv}(c_{1},\dots,c_{d-2}))$ and $\operatorname{vol}_{d-3}(\operatorname{conv}(\pi_{w^{\perp}}(c_{1}),\dots,\pi_{w^{\perp}}(c_{d-2}))$ are proportional, consider the following. Note that $\pi_{w^{\perp}}$ is a fixed projection. Generically we know that $w\notin\operatorname{span}(b_{2}-b_{1},\dots,b_{d-1}-b_{1})$, this is a non-degeneracy condition that holds true with probability $1$. This implies that $\pi_{w^{\perp}}$ is a bijection between the affine hyperplane $\operatorname{affhull}(b_{1},\dots,b_{d-1})$ and its image $\pi_{w^{\perp}}(\operatorname{affhull}(b_{1},\dots,b_{d-1}))$, and also a bijection between the subsets
$\phi^{\perp}\cap\operatorname{affhull}(b_{1},\dots,b_{d-1})\text{~and~}\pi_{w^{\perp}}(\phi^{\perp}\cap\operatorname{affhull}(b_{1},\dots,b_{d-1})).$
This is the space where $c_{1},\dots,c_{d-2}$ live. It follows that $\operatorname{conv}(c_{1},\ldots,c_{d-2})$ has the same dimension as its projection $\operatorname{conv}(\pi_{w^{\perp}}(c_{1}),\dots,\pi_{w^{\perp}}(c_{d-2}))$. The ratio between their volumes depends only on $w$ and $\phi^{\perp}\cap\operatorname{span}(b_{2}-b_{1},\dots,b_{d-1}-b_{1})$. Note that $h=\phi^{\perp}b_{1}$ and so, after fixing $(\alpha,\phi,c_{1},\dots,c_{d-1},b_{d})$ the ratio between the volumes is constant.
Note that the event $0\in\operatorname{conv}(\pi_{w^{\perp}}(b_{i}):i\in[d-1])$ depends only on these variables and not on $h$, and the same is true for $\operatorname{vol}_{d-3}(\operatorname{conv}(\pi_{w^{\perp}}(c_{1}),\dots,\pi_{w^{\perp}}(c_{d-2})))$. We are looking only at the randomness in $h$, and so we can ignore any constant factors in the probability density function. The induced probability density on $h$ is now proportional to
Next, consider the volume term. We can write $\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\dots,b_{d}))$ as a constant depending on $d$ times the absolute value of the determinant of the following $(d-1)\times(d-1)$ matrix
$\begin{bmatrix}(b_{1}-b_{d})^{\top}\\ \vdots\\ (b_{d-1}-b_{d})^{\top}\end{bmatrix}=\begin{bmatrix}(h\phi+c_{1}-b_{d})^{\top}\\ \vdots\\ (h\phi+c_{d-2}-b_{d})^{\top}\\ (-\alpha h\phi+c_{d-1}-b_{d})^{\top}\end{bmatrix}=\begin{bmatrix}(c_{1}-b_{d})^{\top}\\ \vdots\\ (c_{d-2}-b_{d})^{\top}\\ (c_{d-1}-b_{d})^{\top}\end{bmatrix}+h\cdot\begin{bmatrix}1\\ \vdots\\ 1\\ -\alpha\end{bmatrix}\phi^{\top},$
Define
$B:=\begin{bmatrix}(c_{1}-b_{d})^{\top}\\ \vdots\\ (c_{d-2}-b_{d})^{\top}\\ (c_{d-1}-b_{d})^{\top}\end{bmatrix},v:=\begin{bmatrix}1\\ \vdots\\ 1\\ -\alpha\end{bmatrix}$.
Recall that both $B$ and $v$ are fixed and we are only interested in the distribution of $h$. Then by the matrix determinant lemma, we can write the volume as the absolute value of an affine function of $h$ (which is a convex function):
$k(h):=\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\ldots,b_{d}))$ $\propto\left|\det(B+hv\phi^{\top})\right|$
$=\left|\det(B)(1+j\phi^{\top}B^{-1}v)\right|$
To finalize the argument, we write $\operatorname{dist}(\pi_{w^{\perp}}(b_{i}),\operatorname{affhull}(\pi_{w^{\perp}}(b_{j}):j\in[d-1],j\neq i))=|(1+\alpha)h|$. It follows that the signed distance $(1+\alpha)h$ has a probability density function proportional to the product of a $dL$-log-Lipschitz function and a convex function. The result follows from Lemma 5.25 by plugging in the signed distance $(1+\alpha)h$, $K=dL$, and $\varepsilon=\frac{1}{160d^{2}L}$.
5.8 Combining Together and Proof of Lemma 5.2
@@
In this section, we combine the deterministic argument in Lemma 5.6 with the probabilistic arguments in Lemma 5.10 and Lemma 5.16. We can finally show the main technical lemma (Lemma 5.2).
Proof 5.28 (Proof of Lemma 5.2).
@@
Without loss of generality, let $I=[d]$ and write $E=E_{I}$. Suppose $p^{\prime}=A_{J}=\operatorname{conv}(a_{j}:j\in J)\cap W$ is the next vertex after the edge $F_{I}\cap W$. Here $J\in\binom{[n]}{d-1}$ and $R_{J}$ is the (d2)(d-2)-dimensional ridge. With probability $1$, the polytope $\operatorname{conv}(a_{1},\dots,a_{n})$ is non-degenerate and $W\cap R^{\prime}$ is a single point for any ridge $R^{\prime}$ of $\operatorname{conv}(a_{1},\dots,a_{n})$ that intersects with $W$. We will show that conditional on $E$, each of the following conditions in the deterministic argument (Lemma 5.6) is satisfied with good probability:
@@
    (Lower bound of $\delta$) $min_{k\in[n]\backslash I}\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})\geq\Omega(\frac{1}{Ld\log n})$;
@@
Note for the last point that Lemma 5.6 only requires this for the set $J$ which indexes the second vertex of $F_{I}\cap W$ in clockwise direction, but we prove it for both of the sets $J$ for which $R_{J}\cap W\neq\emptyset$.
First, we write $D$ as the event that $\forall i,j\in[n]$ for which $\|a_{i}-a_{j}\|\leq 3$. From Lemma 2.10, for any $\sigma\leq\frac{1}{8\sqrt{d\log n}}$, with probability at least $1-\binom{n}{d}^{-1}$, we have $\max_{i\in[n]}\|\hat{a}_{i}\|\leq 1+4\sigma\sqrt{d\log n}\leq\frac{3}{2}$, i.e., $\Pr[D]\geq 1-\binom{n}{d}^{-1}$. Using the assumption that $\Pr[E_{I}]\geq 10\binom{n}{d}^{-1}$, we have
@@
Next, we consider $\delta:=\operatorname{dist}(\operatorname{affhull}(a_{1},\dots,a_{d}),\{a_{d+1},\dots,a_{n}\})$. Using Lemma 5.10, we have $\Pr[\delta\geq\frac{1}{10e^{3}Ld\log n}\mid E]\geq 0.72$.
Finally, we consider $r:=\max_{J}\operatorname{dist}(\operatorname{affhull}(a_{1},\ldots,a_{d})\cap W,\partial R_{J})$ subject to all $J\in\binom{I}{d-1}$ such that $A_{J}$ happens (in other words, $R_{J}=\operatorname{conv}(a_{j}:j\in J)$ is a ridge of $F_{I}$ such that $R_{J}\cap W\neq\emptyset$). By a union bound,
@@
From Lemma 5.16, for each $J\in\binom{I}{d-1}$, we know that
@@
$\geq$ $1-0.1\cdot 2-2\cdot\Pr[\neg D\mid E]\geq 0.6.$
Therefore, by a union bound, the three conditions hold with probability at least $1-(1-0.9)-(1-0.72)-(1-0.6)\geq 0.1$, and the lemma directly follows from Lemma 5.6.
@@ 6 Smoothed Complexity Lower Bound
In this section, we present the lower bound of the smoothed complexity of the shadow vertex simplex method by studying the intersection of the smoothed polar polytope $\operatorname{conv}(a_{1},\ldots,a_{n})\subseteq\mathbb{R}^{d}$ (where each $a_{i}$ is a Gaussian perturbation from a $\bar{a}_{i}$), and the two-dimensional shadow plane $W\subseteq\mathbb{R}^{d}$. Our main result is as follows:
Theorem 6.1.
For any d>5,n=4d13d>5,n=4d-13, there exists a two-dimensional linear subspace $W\subseteq\mathbb{R}^{d}$ and vectors $\bar{a}_{1},\ldots,\bar{a}_{n}\in\mathbb{R}^{d}$, $\max_{i\in n}\|\bar{a}_{i}\|_{1}\leq 1$ such that the following holds. Let $a_{1},\ldots,a_{n}$ be independent Gaussian random variables where each $a_{i}\sim\mathcal{N}_{d}(\bar{a}_{i},\sigma^{2}I),\sigma\leq\frac{1}{7200d\sqrt{\log n}}$. Then with probability at least $1-\binom{n}{d}^{-1}$ we have
@@
Theorem 6.1 is a direct consequence of the next theorem, which is a lower bound for adversarial perturbations of bounded magnitude:
Theorem 6.2.
For any d>5,n=4d13d>5,n=4d-13, there exists a two-dimensional linear subspace $W\subseteq\mathbb{R}^{d}$ and vectors $\bar{a}_{1},\ldots,\bar{a}_{n}\in\mathbb{R}^{d}$, $\max_{i\in n}\|\bar{a}_{i}\|_{1}\leq 1$ such that the following holds. For any $\varepsilon<\frac{1}{180}$, if $a_{1},\ldots,a_{n}\in\mathbb{R}^{d}$ satisfy $\|a_{i}-\bar{a}_{i}\|_{1}\leq\varepsilon$ for all $i\in[n]$ then we have
@@
The rest of this section is organized as follows. In Section 6.1, we construct an auxiliary polytope $P\subseteq\mathbb{R}^{d}$ and a two-dimensional shadow plane $W$. In Section 6.2, we show that the projection $\pi_{W}(P)$ approximates the unit disk $\mathbb{B}_{2}^{2}$. In Section 6.3, we analyse the largest $\ell_{\infty}$-ball contained in $P$ and the smallest $\ell_{\infty}$-ball containing $P$. Section 6.4 investigates the polar polytope $Q=(P-x)^{\circ}$ of a shift of $P$, such that we may choose $\bar{a}_{1},\dots,\bar{a}_{n}$ to satisfy $Q=\operatorname{conv}(\bar{a}_{1},\dots,\bar{a}_{n})$. The largest contained $\ell_{1}$-ball in $Q$ is derived from the smallest $\ell_{\infty}$-ball containing $P$, and the smallest $\ell_{1}$-ball containing $Q$ is derived from the largest $\ell_{\infty}$-ball contained in $P$. Similarly, the section $Q\cap W$ approximates a circular disk because the projection $\pi_{W}(P)$ approximates a circular disk. Finally, Section 6.5 shows that the bounded ratio between the inner and outer radii implies that any sufficiently small perturbation $\tilde{Q}$ still has $\tilde{Q}\cap W$ approximate the unit disk $\mathbb{B}_{2}^{2}$ well and uses this to prove Theorem 6.1.
6.1 Construction of the Auxiliary Polytope
In this subsection, we construct the auxiliary polytope $P$ and the two-dimensional plane $W$. For $k\in\mathbb{N}$, we construct a $(k+5)$-dimensional polytope. We will use the following vectors in the definition:
@@ 6.1 Construction of the Auxiliary Polytope
With these definitions in mind, construct an auxiliary $P^{\prime}\subseteq\mathbb{R}^{3k+5}$ as the set of points $(x,y,p_{0},\dots,p_{k},t,s)$ where $x,y\in\mathbb{R},p_{0},p_{1},\ldots,p_{k}\in\mathbb{R}^{2},t\in\mathbb{R}^{k}$ and $s\in\mathbb{R}$ satisfy the following system of linear inequalities:
@@
$\mathbf{0}_{k}\leq t$ $\leq\mathbf{1}_{k}$ (28)
@@
We choose the plane $W$ to be the one that is spanned by the unit vectors in the $x$ and $y$ directions.
An illustration of the vertices of the projected polytope $\pi_{W}(P)$ can be found in Section 6.1 for $k=4$. Note that the figure appears to depict a regular polygon with $2^{k+1}$ vertices. Our construction of $P^{\prime}$ is similar to those of [8, 28]. The primary difference lies in the addition of the variables $t$ and $s$ in (26). This change is made to ensure that the projected polytope $P$ has its largest contained and smallest containing $\ell_{\infty}$-ball be of similar sizes.
{sidefigure}[Uncaptioned image]
Vertices of the projected auxiliary polytope $\pi_{W}(P)$ (see (30)) without perturbation for $k=4$.
6.2 Projected Auxiliary Polytope Approximates Two-Dimensional Unit Disk